Circular No. 7327

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1997 K2 T. Mäkinen, Finnish Meteorological Institute, Helsinki; and J.-L.

Bertaux, Service d’Aeronomie, Verrières-le-Buisson, report for the SWAN team the probable discovery of a comet from its hydrogen Lyα emission appearing on full-sky images (wavelength range 10–180 nm) taken from the SOHO spacecraft during May–July 1997:

1997 UT  α_{2000}  δ_{2000}
May 20.884  0^{h}51.^{m}1  −22^{o}16′

The resolution is 1 square degree. Further positional and orbital information is given on MPEC 1999-X07. The discovery was made by Mäkinen in the course of a complete survey of visible comets in the SWAN images during Jan. 1996–June 1998. The appearance was comparable to that of comet C/1997 O1 (Tilbrook) in SWAN images during 1997 June 17–Sept. 6, suggesting that m_{1} was roughly 10–12. Peak Lyα intensity values (in Rayleighs) for C/1997 K2 were: June 5.0 UT, 10.4; 7.4, 12.9; 13.1, 12.3; 14.8, 14.7; 24.5, 17.9; 26.6, 17.9; 28.7, 17.4. The object increased in extent from 2^{o} × 2^{o} to 6^{o} × 4^{o} during this time.

V1494 AQUILAE W. Liller, Viña del Mar, Chile, reports that a CCD spectrogram ob-

tained of V1494 Aql on Dec. 3.0346 UT with an objective prism shows Hα in emission at a level 31 percent above the neighboring continuum; Hβ is faintly visible.

A. Frasca and S. Messina, Osservatorio Astrofisico di Catania, report Strömgren photometry of V1494 Aql at Serra La Nave on Mount Etna with the 0.91-m telescope on Dec. 3.712 UT (over ∼ 0.5 hr): V = 4.025 ± 0.037, b−y = +0.536 ± 0.014, m_{1} = 0.054 ± 0.019, c_{1} = 0.978 ± 0.037. Variations of ∼ 0.06–0.08 mag were observed in all filters.

Visual magnitude estimates: Dec. 3.702 UT, 4.1 (R. J. Bouma, Groningen, The Netherlands); 4.378, 4.6 (S. Yoshida, Ibaraki, Japan); 4.670, 4.7 (K. Sárneczky, Budapest, Hungary); 4.774, 5.2 (C. Vitorino, Cabo da Roca, Portugal); 4.776, 4.8 (A. Pereira, Cabo da Roca, Portugal); 4.778, 5.0 (A. P. da Silva Correia and J. Rodrigues Ribeiro, Esposende, Portugal); 4.850, 5.2 (Vitorino); 4.852, 5.1 (Pereira).

1999 December 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green